(D2-D'2-3D+3D')z=xy+7


1
Expert's answer
2021-05-28T10:01:15-0400

Solution:-

"(D^2-D'^2-3D+3D')z=xy+7"

"(D-D')(D+D'-3)z=xy+7"

"m_1=1 \\
m_2=-1 \\
c_1=0 \\
c_2=3"


"c.f=e^{c_1x}f_1(y+m_1x)+e^{cx}f_2(y+m_2x)"

"c.f=e^{0x}f_1(y+x)+e^{3x}f_2(y-x)"

"c.f=f_1(y+x)+e^{3x}f_2(y-x)"


PI="\frac{xy+7}{(D-D')(D+D'-3)}"

"D'=-1\\
D=1"

"PI = \frac{xy+7}{-6}"


answer = C.F + P.I

="f_1(y+x)+e^{3x}f_2(y-x) - (\frac{xy+7}{6})"